KiWiPoP1:

what is the sum of √(12) + 5 √(3)

Laylalyssa:

@jhonyy9 please

heavystepper:

@darkknight

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @heavystepper @darkknight \(\color{#0cbb34}{\text{End of Quote}}\) he's not even online-

heavystepper:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa \(\color{#0cbb34}{\text{Originally Posted by}}\) @heavystepper @darkknight \(\color{#0cbb34}{\text{End of Quote}}\) he's not even online- \(\color{#0cbb34}{\text{End of Quote}}\) ...

heavystepper:

@jhonyy9

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @heavystepper @jhonyy9 \(\color{#0cbb34}{\text{End of Quote}}\) I already tagged him

KiWiPoP1:

welp, i got it

heavystepper:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KiWiPoP1 welp, i got it \(\color{#0cbb34}{\text{End of Quote}}\) ok

KiWiPoP1:

ty ty

Trios:

I used a site called mathpapa and got 7√3

KiWiPoP1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Trios I used a site called mathpapa and got 7√3 \(\color{#0cbb34}{\text{End of Quote}}\) yeah i got that

KiWiPoP1:

imma close this now 🌚

Trios:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KiWiPoP1 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Trios I used a site called mathpapa and got 7√3 \(\color{#0cbb34}{\text{End of Quote}}\) yeah i got that \(\color{#0cbb34}{\text{End of Quote}}\) Ah well I’m late lmao

Trios:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @KiWiPoP1 imma close this now 🌚 \(\color{#0cbb34}{\text{End of Quote}}\) Yea 🌝

jhonyy9:

sqrt 12 +5sqrt3 = sqrt(4*3) +5sqrt3 = 2sqrt3 +5sqrt3 = 7sqrt3

jhonyy9:

@Laylalyssa

Laylalyssa:

oh is that it..

jhonyy9:

yeah so easy to solve it

Laylalyssa:

I thought √(12) + 5 √(3) 5+3√3

jhonyy9:

how you get this ? step by step please ?

Laylalyssa:

√12+5+√3 √12= 2√3 √12+5=5+2√3 5+3√3 no?

jhonyy9:

this 5sqrt3 is 5*sqrt3 not plus

Laylalyssa:

oh

jhonyy9:

5sqrt3 = sqrt(25*3) = sqrt75

jhonyy9:

ok ?

Laylalyssa:

ok ok

jhonyy9:

questions ?

Laylalyssa:

no ty

jhonyy9:

yw bye bye